Carla and David were candidates for a student government office. A total of 8532 votes were cast for these two candidates, and Carla won by getting 692 more votes than David. How many votes did each candidate get?
Answer by richwmiller(17219)

c=d+692
c-d=692
c+d=8532
add the 2nd and third
2c=692+8532